# Nationality and Region Prediction from Names: A Comparative Study of Neural Models and Large Language Models

### v1

@@ 4.4 Evaluation on Robustness to Label Frequency
SVM 46.97% 48.94% 46.85% +0.12% 0.3%
fastText 28.14% 26.10% 20.14% +8.00% 28.4%
CNN 43.94% 42.50% 38.29% +5.65% 12.9%
BiLSTM 40.72% 41.21% 35.85% +4.87% 12.0%
CANINE 47.05% 44.55% 37.89% +9.16% 19.5%
XLM-RoBERTa 48.07% 44.85% 37.22% +10.85% 22.6%
@@
Zero-shot 79.39% 74.73% 72.49% +6.90% 8.7%
Few-shot 74.91% 71.66% 70.77% +4.15% 5.5%
Chain-of-Thought 74.70% 68.37% 65.96% +8.74% 11.7%
Self-Consistency 79.05% 75.23% 72.76% +6.29% 8.0%
Least-to-Most 66.25% 59.05% 58.96% +7.29% 11.0%
Self-Reflection 80.83% 76.67% 73.43% +7.40% 9.2%
Among neural models, pre-trained models exhibited the most pronounced frequency dependence. XLM-RoBERTa showed $\Delta$(H-T) of +10.85% and CANINE showed +9.16%, with performance in the Tail group decreasing by approximately 20% compared to the Head group. This suggests that pre-trained models are influenced by the frequency of languages and nationalities in the pre-training corpus. Since pre-training corpora contain more text in English and European languages, name patterns from these regions may be better learned.
In contrast, SVM showed $\Delta$(H-T) of +0.12%, nearly zero, demonstrating extremely high robustness. Since SVM is based on character n-gram pattern matching, it directly reflects the frequency distribution of training data, making it less prone to bias toward specific nationalities. fastText showed $\Delta$(H-T) of +8.00%, but its Drop% was the highest at 28.4%, indicating that its already low performance deteriorated further in the Tail group.
For LLMs, $\Delta$(H-T) ranged from +4% to +9% across all methods. Few-shot was the most robust, with $\Delta$(H-T) of +4.15% and Drop% of only 5.5%. Since Few-shot selects examples from each region, examples from regions containing low-frequency nationalities may have had an effect of mitigating bias. On the other hand, Chain-of-Thought showed a relatively high $\Delta$(H-T) of +8.74%, suggesting that biased judgments toward high-frequency nationalities are more likely to occur during the reasoning process.
Focusing on Macro-F1, the tendency of frequency dependence became more pronounced. LLM Macro-F1 ranged from 0.35 to 0.44 in the Head group but decreased to 0.26 to 0.31 in the Tail group. For neural models, it ranged from 0.13 to 0.19 in the Head group and 0.07 to 0.18 in the Tail group. Since Macro-F1 weights each class equally, performance degradation for individual nationalities in the Tail group is directly reflected. While LLMs show high robustness in Accuracy, performance degradation for low-frequency nationalities was clearly observed in Macro-F1.
@@ 4.5 Error Analysis
First, we analyzed whether errors in nationality prediction are also errors at the region level, or become correct at the region level. Table 9 shows the results of region-level accuracy analysis. Even when LLM made incorrect nationality predictions, it predicted the correct region in 86.2% of cases. In contrast, XLM-RoBERTa’s region accuracy remained at 66.8%. This difference indicates that LLM possesses extensive knowledge about the relationships between names and regions, capturing the general direction even when making fine-grained errors.
@@
Nationality Correct 5,703 (75.7%) 3,292 (43.7%)
Nationality Wrong, Region Correct 788 (10.5%) 1,743 (23.1%)
Nationality Wrong, Region Wrong 1,043 (13.8%) 2,499 (33.2%)
Region Accuracy (incl. Nationality Correct) 86.2% 66.8%
@@
Table 10 shows the top-15 confusion pairs and region agreement rates for both models. 66.7% of LLM’s confusion pairs are errors within the same region. The most frequent confusion is English → British (67 cases), which is an error between extremely close nationalities within the same Western Europe region. Similarly, confusions between linguistically and culturally proximate nationalities such as Tamil → Indian, Welsh → British, and Uruguayan → Argentine are common. These errors are considered cases where distinction is inherently difficult due to extremely similar name patterns.
Table 10: Top-15 confusion pairs.
@@
Region Agreement 66.7% Region Agreement 60.0%
For XLM-RoBERTa’s confusion pairs, the region agreement rate is 60.0%, lower than LLM. A distinctive characteristic is the tendency for South American nationalities (Peruvian, Ecuadorian, Colombian, Guatemalan, Cuban) to be misclassified as Mexican. This is likely due to the high frequency of Mexican in the training data, resulting in prediction bias toward Mexican for Spanish-speaking names. Cross-regional errors such as Brazilian → Portuguese are also observed, indicating misclassification based on linguistic similarity (Portuguese language).

### v2

@@ 4.4 Evaluation on Robustness to Label Frequency
SVM 0.470 0.489 0.469 +0.001 0.3%
fastText 0.281 0.261 0.201 +0.080 28.4%
CNN 0.439 0.425 0.383 +0.057 12.9%
BiLSTM 0.407 0.412 0.359 +0.049 12.0%
CANINE 0.471 0.446 0.379 +0.092 19.5%
XLM-RoBERTa 0.481 0.449 0.372 +0.109 22.6%
@@
Zero-shot 0.794 0.747 0.725 +0.069 8.7%
Few-shot 0.749 0.717 0.708 +0.042 5.5%
Chain-of-Thought 0.747 0.684 0.660 +0.087 11.7%
Self-Consistency 0.791 0.752 0.728 +0.063 8.0%
Least-to-Most 0.663 0.591 0.590 +0.073 11.0%
Self-Reflection 0.808 0.767 0.734 +0.074 9.2%
Among neural models, pre-trained models exhibited the most pronounced frequency dependence. XLM-RoBERTa showed $\Delta$(H-T) of 0.109 and CANINE showed 0.092, with performance in the Tail group decreasing by approximately 20% compared to the Head group. This suggests that pre-trained models are influenced by the frequency of languages and nationalities in the pre-training corpus. Since pre-training corpora contain more text in English and European languages, name patterns from these regions may be better learned.
In contrast, SVM showed $\Delta$(H-T) of only 0.001, demonstrating extremely high robustness. Since SVM is based on character n-gram pattern matching, it directly reflects the frequency distribution of training data, making it less prone to bias toward specific nationalities. fastText showed $\Delta$(H-T) of 0.080, but its Drop% was the highest at 28.4%, indicating that its already low performance deteriorated further in the Tail group.
For LLMs, $\Delta$(H-T) ranged from 0.042 to 0.087 across all methods. Few-shot was the most robust, with $\Delta$(H-T) of 0.042 and Drop% of only 5.5%. Since Few-shot selects examples from each region, examples from regions containing low-frequency nationalities may have had an effect of mitigating bias. On the other hand, Chain-of-Thought showed a relatively high $\Delta$(H-T) of 0.087, suggesting that biased judgments toward high-frequency nationalities are more likely to occur during the reasoning process.
Focusing on Macro-F1, the tendency of frequency dependence became more pronounced. LLM Macro-F1 ranged from approximately 0.350 to 0.440 in the Head group but decreased to 0.260 to 0.310 in the Tail group. For neural models, it ranged from approximately 0.130 to 0.190 in the Head group and 0.070 to 0.180 in the Tail group. Since Macro-F1 weights each class equally, performance degradation for individual nationalities in the Tail group is directly reflected. While LLMs show high robustness in Accuracy, performance degradation for low-frequency nationalities was clearly observed in Macro-F1.
@@ 4.5 Error Analysis
First, we analyzed whether errors in nationality prediction are also errors at the region level, or become correct at the region level. Table 9 shows the results of region-level accuracy analysis. Even when LLM made incorrect nationality predictions, it predicted the correct region in 0.862 of cases. In contrast, XLM-RoBERTa’s region accuracy remained at 0.668. This difference indicates that LLM possesses extensive knowledge about the relationships between names and regions, capturing the general direction even when making fine-grained errors.
@@
Nationality Correct 5,703 (0.757) 3,292 (0.437)
Nationality Wrong, Region Correct 788 (0.105) 1,743 (0.231)
Nationality Wrong, Region Wrong 1,043 (0.138) 2,499 (0.332)
Region Accuracy (incl. Nationality Correct) 0.862 0.668
@@
Table 10 shows the top-10 confusion pairs and region agreement rates for both models. 0.667 of LLM’s confusion pairs are errors within the same region. The most frequent confusion is English → British (67 cases), which is an error between extremely close nationalities within the same Western Europe region. Similarly, confusions between linguistically and culturally proximate nationalities such as Tamil → Indian, Welsh → British, and Uruguayan → Argentine are common. These errors are considered cases where distinction is inherently difficult due to extremely similar name patterns.
Table 10: Top-10 confusion pairs.
@@
Region Agreement 0.667 Region Agreement 0.600
For XLM-RoBERTa’s confusion pairs, the region agreement rate is 0.600, lower than LLM. A distinctive characteristic is the tendency for South American nationalities (Peruvian, Ecuadorian, Colombian, Guatemalan, Cuban) to be misclassified as Mexican. This is likely due to the high frequency of Mexican in the training data, resulting in prediction bias toward Mexican for Spanish-speaking names. Cross-regional errors such as Brazilian → Portuguese are also observed, indicating misclassification based on linguistic similarity (Portuguese language).